OpenStudy (anonymous):

Okay i have no idea how to do this, and nobody at my house knows so please answer :) "Write an equation of the line parallel to given line containing C." 12. C(6,-2);y= -3/2x+6 Then another ones, "Write an equation of line perpendicular to the given line contains P." 1.P(6,6);y=2/3x 2.P(4,4);y=-2x-8

OpenStudy (anonymous):

Helppp :(

OpenStudy (anonymous):

slope of \(y=-\frac{3}{2}x+6\) is \(-\frac{3}{2}\) your job is to find the equation of the line with slope \(-\frac{3}{2}\) through the point \((6,-2)\) the point-slope formula \[y=y_1=m(x-x_1)\] will give it to you use \(m=\frac{3}{x},x_1=6,y_1=-2\)

OpenStudy (anonymous):

oops i meant \(m=-\frac{3}{2}\)

OpenStudy (anonymous):

The equation that you are looking for y=mx+b Parallel with line y=-3/2x+6 Therefore m=-3/2 Now your line looks like y= -3/2x +b. this line passes the point C(6,-2)next step we have -2=-3/2(6)+b -2=-9+b -2+9=b b=7 Final you line equaltion is y=-3/2+7 this line parallel iwth the line y=-3/2x+6 and also passes the point C(6,-2)

OpenStudy (anonymous):

so u will get y=-3/2x+7 solving using the point slope formula

OpenStudy (anonymous):

either way will work

OpenStudy (anonymous):

How do i know if it's parrell?

OpenStudy (anonymous):

Because parallel lines have same slopes, perpendicular negative reciprocals, and intersecting lines different slopes

OpenStudy (anonymous):

Thank you guys! <3

OpenStudy (anonymous):

welcomee